OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, -1). Answer y =1/3 x - 3 y = 1/3 x + 17 y = -3x - 3 y = -3x + 17

OpenStudy (aroub):

y=mx+b , m=1/3 ; (6,-1) -1=1/3(6)+b -1=2+b b=-3 => y=1/3x-3